# Optimal control of discrete-time nonlinear systems with application to automatic guided vehicles

### v2

@@ I Introduction
In this paper, the OCP for discrete-time nonlinear systems is considered. Unlike previous research that relies on existing optimization algorithms [13, 14, 15, 16, 17], this work presents a novel optimization approach from the perspective of optimal control, featuring a superlinear convergence property. Following the work of [18] and [19], this study specifically addresses dynamic problems, with a particular focus on the challenges associated with effectively computing the gradient and hessian matrix. To tackle these challenges, a new approach based on the maximum principle is introduced for the first time, transforming the solution process into solving two sets of FBDEs. This provides a unified computational approach for developing second-order optimal control algorithms. The contributions of this paper can be summarized as follows:
@@ III-A Optimal Control Design
Although FBDEs (6)-(8) can be numerically solved using the existing MSA, its convergence rate is relatively slow. To improve computational efficiency, an iterative numerical algorithm has been developed by combining the Taylor expansion with mathematical induction in [18, 19], as outlined below.
$z_{i+1}$ $=z_{i}-g_{i},$ (9)
$g_{i}$ $=(R+\nabla^{2}J(x_{0},z_{i}))^{-1}(\nabla J(x_{0},z_{i})+Rg_{i-1}),$ (10)
$g_{0}$ $=(R+\nabla^{2}J(x_{0},z_{i}))^{-1}\nabla J(x_{0},z_{i}),$ (11)
where $i=0,1,...,M-1$.
@@
In the iterative algorithm (9)-(11), the update step $g_{i}$ requires inner loop iterations starting from $g_{0}$. It should be noted that the algorithm faces considerable challenges in handling dynamic problem, i.e., the OCP (1)-(2), due to the computational complexity involved in evaluating $\nabla J$ and $\nabla^{2}J$. The following subsections provide the explicit iterative solution algorithms. For details on the convergence analysis and the superlinear convergence rate of the algorithm (9)-(11), please refer to [18, 19].

### v3

@@ I Introduction
In this paper, the OCP for discrete-time nonlinear systems is considered. Unlike previous research that relies on existing optimization algorithms [13, 14, 15, 16, 17], this work presents a novel optimization approach from the perspective of optimal control, featuring a superlinear convergence property. Following the work of [18], this study specifically addresses dynamic problems, with a particular focus on the challenges associated with effectively computing the gradient and hessian matrix. To tackle these challenges, a new approach based on the maximum principle is introduced for the first time, transforming the solution process into solving two sets of FBDEs. This provides a unified computational approach for developing second-order optimal control algorithms. The contributions of this paper can be summarized as follows:
@@ III-A Optimal Control Design
Although FBDEs (6)-(8) can be numerically solved using the existing MSA, its convergence rate is relatively slow. To improve computational efficiency, an iterative numerical algorithm has been developed by combining the Taylor expansion with mathematical induction in [18], as outlined below.
$z_{i+1}$ $=z_{i}-g_{i}(z_{i}),$ (9)
$g_{j}(z_{i})$ $=(R+\nabla^{2}J(x_{0},z_{i}))^{-1}(\nabla J(x_{0},z_{i})+Rg_{j-1}(z_{i})),$ (10)
$g_{0}(z_{i})$ $=(R+\nabla^{2}J(x_{0},z_{i}))^{-1}\nabla J(x_{0},z_{i}),$ (11)
where $i=0,1,...,M-1$ and $j=1,...,i$.
@@
In the iterative algorithm (9)-(11), the update step $g_{i}$ requires inner loop iterations starting from $g_{0}$. It should be noted that the algorithm faces considerable challenges in handling dynamic problem, i.e., the OCP (1)-(2), due to the computational complexity involved in evaluating $\nabla J$ and $\nabla^{2}J$. The following subsections provide the explicit iterative solution algorithms. For details on the convergence analysis and the superlinear convergence rate of the algorithm (9)-(11), please refer to [18].